Two kids are on a seesaw. The one on the left has a mass of 75 kg and is sitting 1.5 m from the pivot point. The one on the right has a mass of 60 kg and is sitting 1.8 m from the pivot point. What is the net torque on the system?

a) 4.5 Nm ccw

b) 44.1 Nm ccw

c) 2160.1 Nm cw

d) 147 Nm cw
Answer #1

Gravitational acceleration = g = 9.8 m/s2

Mass of the kid on the left = m1 = 75 kg

Mass of the kid on the right = m2 = 60 kg

Distance of the kid on the left from the pivot = R1 = 1.5 m

Distance of the kid on the right from the pivot = R2 = 1.8 m

The kid on the left will create a torque in the counter clockwise direction while the kid on the right will create a torque in the clockwise direction.

Torque due to the left side kid = T1

T1 = m1gR1

T1 = (75)(9.8)(1.5)

T1 = 1102.5 Nm ccw

Torque due to the right side kid = T2

T2 = m2gR2

T2 = (60)(9.8)(1.8)

T2 = 1058.4 Nm cw

Net torque on the system = T

T = T1 - T2

T = 1102.5 - 1058.4

T = 44.1 Nm ccw

Option (b) is correct